OpenStudy (anonymous):

In Clark Middle School, there are 60 students in seventh grade. If 25 of these students take art only, 18 take music only, and 9 do not take either art or music, how many taky both art and music? answer: ? students

OpenStudy (anonymous):

i think the answer is 8 beacuse if you add the 25 + 18= 43 43=9=52 52-60=8 8 equals the students that takes voth art music class

OpenStudy (anonymous):

sorry i ment 43+9=52

OpenStudy (anonymous):

do you understand